Given two consecutive integers. The sum of 3 times the 1st integer and four times the 2nd integer is 95. Find the integers.
Answer by jim_thompson5910(35256)

x and y are the integers

"Given two consecutive integers" ----> y = x+1

"The sum of 3 times the 1st integer and four times the 2nd integer is 95" ---> 3x+4y = 95

----------------------------------------------------------------------


3x+4y = 95

3x+4(x+1) = 95 ... replace y with x+1

3x + 4x + 4 = 95

7x + 4 = 95


I'll let you finish up.